Biographical Sketch

Biographical Sketch : John B. Little

Department of Mathematics and Computer Science College of the Holy Cross Worcester, MA 01610, USA

mathcs.holycross.edu/~little/homepage.html

Phone: (508) 793-2274 FAX: (508) 793-3040

                                          little@mathcs.holycross.edu

Professional Preparation Haverford College, AB in Mathematics, 1976 Harvard University, visiting scholar, 1979-1980 Yale University, Ph.D. in Mathematics, 1980

Appointments Professor, College of the Holy Cross, Since September 2003. Mathematical Sciences Research Institute, Berkeley, Member January – June 2003 Associate Professor, College of the Holy Cross, September 1986 – August 2003. Visiting Fellow, Cornell University Mathematical Sciences Institute, September 1993 – June 1994. Assistant Professor, College of the Holy Cross, September 1980 – August 1986.

Publications most related to project (1) Remarks on generalized toric codes, arXiv:1107.4530, submitted.

(2) Toric Varieties, with D. Cox and H. Schenck. Book, xxiv + 841 pages, Graduate Studies in Mathematics 124, AMS, 2011.

(3) On toric codes and multivariate Vandermonde matrices, with R. Schwarz. Applicable Algebra in Engineering, Communications, and Computing 18 (2007), 349-367.

(4) Toric surface codes and Minkowski sums, with H. Schenck. SIAM Journal on Discrete

Mathematics 20 (2006), 999-1014. Other Significant Publications

(1) Using Algebraic Geometry, with D. Cox and D. O’Shea. Book, xi + 572 pages, Graduate Texts in Mathematics 185, Springer, 1998, second edition, 2005.

(2) Ideals, Varieties, and Algorithms, with D. Cox and D. O’Shea. Book, xv + 551 pages, Undergraduate Texts in Mathematics, Springer 1992, second edition, 1996, third edition 2007.

Synergistic Activities

(1) Seminar Leader, PURE Math 2012 REU-type program, University of Hawai’i at Hilo, 11 June – 10 August, 2012.

(2) Seminar Leader, MSRI-UP 2009 REU, MSRI, Berkeley, 15 June – 24 July, 2009.

(3) Seminar Leader, PREMUR REU, University of Puerto Rico at Mayaguez, 11 June – 27 July, 2007.

(4) Seminar Leader, SIMU REU-type program, University of Puerto Rico at Humacao, summers 1998,1999,2001.

Collaborators and Other Affiliations D. Cox (Amherst College) J. Hachmeister (undergraduate student, U. Hawai’i at Hilo) J. McGhee (undergraduate student, Loyola Marymount University) D. O’Shea (Mount Holyoke College, New College of Florida) R. Pelayo (U. Hawai’i at Hilo) S. Sasarita (undergraduate student, University of Arizona) H. Schenck (U.Illinois Urbana-Champaign) R. Schwarz (SUNY New Paltz)

Thesis Advisor: Bernard Saint-Donat